OpenStudy (darrenmadx):

Solve the equation. y + (–6) = –9 A. –15 B. –3 C. 54 D. 3

OpenStudy (anonymous):

y + (–6) = –9 or, y-6=-9 or,y=-9+6 or,y=-3

OpenStudy (anonymous):

so , it's B

OpenStudy (anonymous):

@DarrenMadx

OpenStudy (darrenmadx):

oh thanks for the help

OpenStudy (anonymous):

np:)